OpenStudy (nannie):

ANYBODY????? HELP! If 1570536900*3962.946591^-2=100.0026954 what does 1570536900*3959.15156^-2=??? I am no able to get ^-2. KEEP GETTING LOST!

OpenStudy (amistre64):

calculator if need be, this is why we have computers and such

OpenStudy (anonymous):

i don't see a quick gimmick here you have \[\frac{a}{b^2}=c\] and you want \[\frac{a}{(b-3.795031)^2}\]

OpenStudy (saifoo.khan):

Heyyyy!

OpenStudy (saifoo.khan):

Still stuck?

OpenStudy (nannie):

Had Dr. Appointment. Yep , still stuck. my calculator doesn't get the ^-2

OpenStudy (nannie):

I want to see how the results come from.